An urn contains 12 white, 5 yellow and 13 black marbles, a marble is chosen at randok from the urn, and it is noted that its not one of black marbles. What is the sample space in view of this knowledge? Whats the probability thats yellow?


1
Expert's answer
2021-03-18T15:45:17-0400
"S=\{12white, 5yellow\}"

"P(yellow)=\dfrac{5}{5+12}=\dfrac{5}{17}"